if my mother is currently three times my age and in 15 years she will be twice my age how old am i now?
Answer by stanbon(75887)

if my mother is currently three times my age and in 15 years she will be twice my age how old am i now?
----
Equations:
m = 3x
(m+15) = 2(x+15)
----
Substitute for "m" and solve for "x":
3x+15 = 2x+30
x = 15 ("My" age now)
===========================
Cheers,
Stan H.